# Search for the radiative transition $\chi_{c1}(3872)\to\gamma\psi_{2}(3823)$

### v1

@@ Abstract
Using 9.0 $\rm fb^{-1}$ of $e^{+}e^{-}$ collision data collected at center-of-mass energies from 4.178 to 4.278 GeV with the BESIII detector at the BEPCII collider, we perform the first search for the radiative transition $\chi_{c1}(3872)\to\gamma\psi_{2}(3823)$. No $\chi_{c1}(3872)\to\gamma\psi_{2}(3823)$ signal is observed. The upper limit on the ratio of branching fractions $\mathcal{B}(\chi_{c1}(3872)\to\gamma\psi_{2}(3823),\psi_{2}(3823)\to\gamma\chi_{c1})/\mathcal{B}(\chi_{c1}(3872)\to\pi^{+}\pi^{-}J/\psi)$ is set as 0.075 at the 90% confidence level. Our result contradicts theoretical predictions under the assumption that the $\chi_{c1}(3872)$ is the pure charmonium state $\chi_{c1}(2P)$.
@@ I Introduction
The measurement of new $\chi_{c1}(3872)$ decay modes can help to improve our understanding of its internal structure. Ref. [19] extracted the absolute branching fractions of the known $\chi_{c1}(3872)$ decays by performing a global fit of the absolute branching fraction of the $B^{+}\to\chi_{c1}(3872)K^{+}$ channel measured by BaBar [18] together with information from other experiments. The fraction of $\chi_{c1}(3872)$ decays not observed in experiments is estimated to be $31.9^{+18.1}_{-31.5}$%. If the $\chi_{c1}(3872)$ contains a component of the excited spin-triplet state $\chi_{c1}(2P)$, then the radiative decay $\chi_{c1}(3872)\to\gamma\psi_{2}(3823)$ could happen naturally via a E1 transition [20], where the $\psi_{2}(3823)$ is considered as the $1^{3}D_{2}$ charmonium state. The BESIII experiment has reported the observation of $e^{+}e^{-}\to\gamma\chi_{c1}(3872)$ at center-of-mass energies $\sqrt{s}=4.178-4.278$ GeV [21, 22]. Using the $\chi_{c1}(3872)$ signal produced in these data samples, we search for the radiative transition $\chi_{c1}(3872)\to\gamma\psi_{2}(3823)$, where the $\psi_{2}(3823)$ is reconstructed with the cascade decay $\psi_{2}(3823)\to\gamma\chi_{c1}$, $\chi_{c1}\to\gamma J/\psi$, $J/\psi\to\ell^{+}\ell^{-}$ ($\ell=e,\mu$). The branching fraction ratio of this decay relative to the well-established $\chi_{c1}(3872)\to\pi^{+}\pi^{-}J/\psi$ decay, $\mathcal{R}_{\chi_{c1}(3872)}\equiv\mathcal{B}(\chi_{c1}(3872)\to\gamma\psi_{2}(3823),\psi_{2}(3823)\to\gamma\chi_{c1})/\mathcal{B}(\chi_{c1}(3872)\to\pi^{+}\pi^{-}J/\psi)$, is determined in this work.
@@ III Event Selection and result
A four-constraint (4C) kinematic fit is applied to constrain the total four-momentum of the lepton pair and the four photons to that of the colliding beams, to suppress backgrounds and improve the resolution. For events with more than four photons, the combination with the best fit quality corresponding to the minimum fit chi-square, $\chi^{2}_{\rm 4C}$, is retained. The $J/\psi$ is reconstructed by requiring the invariant mass, $M(\ell\ell)$, of the lepton pair to satisfy $|M(\ell\ell)-m(J/\psi)|<30$ MeV/$c^{2}$, where $m(J/\psi)$ is the nominal $J/\psi$ mass. The dominant background is from the process $e^{+}e^{-}\to\pi^{0}\pi^{0}J/\psi$. After the $J/\psi$ selection, we veto $\pi^{0}$ candidates by requiring that the invariant mass of all photon pairs is more than 15 MeV/$c^{2}$ away from the nominal $\pi^{0}$ mass. After these requirements, a seven-constraint (7C) kinematic fit with an additional three constraints on the masses of $M(\ell\ell)$, $M(\gamma\ell\ell)$, and $M(\gamma\gamma\ell\ell)$ to the nominal masses of $J/\psi$, $\chi_{c1}$, and $\psi_{2}(3823)$, respectively, is applied to distinguish the radiative photon in each cascade decay. The best-fit combination with the minimum chi-square, $\chi^{2}_{\rm 7C}$, is retained; $\chi^{2}_{\rm 7C}<100$ is also required to further suppress the combinatorial backgrounds.
Figure 1 shows the distribution of the invariant mass of the radiative photon and the $\psi_{2}(3823)$, $M(\gamma\psi_{2}(3823))$ for the selected candidates, summed over all the energy points. No signal is observed in the $\chi_{c1}(3872)$ signal region in data. After normalizing the MC samples according to the luminosity and cross section in data, the contributions of the $e^{+}e^{-}\to\pi^{0}\pi^{0}J/\psi$ process and of the other backgrounds, estimated with the inclusive MC sample, are also shown in Fig. 1.
@@
The branching ratio $\mathcal{R}_{\chi_{c1}(3872)}$ is calculated by
@@
where $N_{\rm obs}=0$ is the number of observed events in the $\chi_{c1}(3872)$ signal region [3.855, 3.885] GeV/$c^{2}$ from all data; $N_{\rm obs}^{\rm sdb}=4$ is the number of the observed events in the $\chi_{c1}(3872)$ sideband region [3.840, 3.855] and [3.885, 3.940] GeV/$c^{2}$; $r$, the background scaling factor from the sideband regions to the signal region, is 0.474 based on the inclusive MC sample (taking into account its systematic uncertainty; see Sec. IV); $N_{\pi^{+}\pi^{-}J/\psi}=80.7\pm 9.0$ is taken from the BESIII measurement [10]; the branching fraction $\mathcal{B}(\chi_{c1}\to\gamma J/\psi)=0.343\pm 0.010$ is quoted from the PDG [2]; $\epsilon_{\gamma\psi_{2}(3823)}$ is the efficiency of the signal process determined with the signal MC sample; and $\epsilon_{\pi^{+}\pi^{-}J/\psi}$ is the efficiency of the process $\chi_{c1}(3872)\to\pi^{+}\pi^{-}J/\psi$ [10]. The efficiency ratio $\epsilon_{\gamma\psi_{2}(3823)}/\epsilon_{\pi^{+}\pi^{-}J/\psi}$ at each energy point is shown in Fig. 2, which is almost independent on the center-of-mass energy. The mean value with the error, $\epsilon_{\gamma\psi_{2}(3823)}/\epsilon_{\pi^{+}\pi^{-}J/\psi}=0.433\pm 0.004$, is used to calculate the $\mathcal{R}_{\chi_{c1}(3872)}$ value. The upper limit of $\mathcal{R}_{\chi_{c1}(3872)}$ at the 90% confidence level (C.L.) is computed with the TRolke program implemented in the ROOT framework [33] by assuming the background $N_{\rm obs}^{\rm sdb}$ and the denominator of $\mathcal{R}_{\chi_{c1}(3872)}$ follow Poisson and Gaussian distributions, respectively, where the systematic uncertainties discussed in the following section is taken as the standard deviation of the Gaussian function to be considered in the upper limit. We obtain an upper limit of $\mathcal{R}_{\chi_{c1}(3872)}<0.075$ at the 90% C.L.
@@
Systematic uncertainties on $\mathcal{R}_{\chi_{c1}(3872)}$ arise mainly from the estimations of $r$, $N_{\pi^{+}\pi^{-}J/\psi}$, $\epsilon_{\gamma\psi_{2}(3823)}/\epsilon_{\pi^{+}\pi^{-}J/\psi}$, and $\mathcal{B}(\chi_{c1}\to\gamma J/\psi)$. The background scaling factor $r$ is determined from the inclusive MC samples. We use a 1st or 2nd order polynomial function to fit the $M(\gamma\psi_{2}(3823))$ distribution from the inclusive MC samples; the $r$ value is calculated several times using the parameters from the fit and varying them within $1\sigma$. The value $r=0.474$ is chosen from the obtained values since it provides the most conservative upper limit.

### v2

@@ Abstract
Using 9.0 $\rm fb^{-1}$ of $e^{+}e^{-}$ collision data collected at center-of-mass energies from 4.178 to 4.278 GeV with the BESIII detector at the BEPCII collider, we perform the first search for the radiative transition $\chi_{c1}(3872)\to\gamma\psi_{2}(3823)$. No signal is observed and the upper limit on the ratio of branching fractions $\mathcal{B}(\chi_{c1}(3872)\to\gamma\psi_{2}(3823),\psi_{2}(3823)\to\gamma\chi_{c1})/\mathcal{B}(\chi_{c1}(3872)\to\pi^{+}\pi^{-}J/\psi)$ is set at 0.075 at the 90% confidence level. Our result contradicts theoretical predictions under the assumption that the $\chi_{c1}(3872)$ is the pure charmonium state $\chi_{c1}(2P)$.
@@ I Introduction
Measurements of new $\chi_{c1}(3872)$ decay modes can help to improve our understanding of its internal structure. Ref. [18] extracted the absolute branching fractions of the known $\chi_{c1}(3872)$ decays by performing a global fit of the absolute branching fraction of the $B^{+}\to\chi_{c1}(3872)K^{+}$ channel measured by BaBar [19] together with information from other experiments. The fraction of $\chi_{c1}(3872)$ decays not observed in experiments is estimated to be $31.9^{+18.1}_{-31.5}$%. The work is carried out by assuming the $\chi_{c1}(3872)$ has universal properties in different production and decay mechanisms. Meanwhile, Ref. [20] also reported the branching fractions with consideration of the threshold effect of $D^{*0}\bar{D^{0}}+c.c.$ and a possible bound state below the threshold or a virtual state in $B^{+}\to\chi_{c1}(3872)K^{+}$ decay. If the $\chi_{c1}(3872)$ contains a component of the excited spin-triplet state $\chi_{c1}(2P)$, then the radiative decay $\chi_{c1}(3872)\to\gamma\psi_{2}(3823)$ could happen naturally via a E1 transition [21], where the $\psi_{2}(3823)$ is considered as the $1^{3}D_{2}$ charmonium state. The BESIII experiment has reported the observation of $e^{+}e^{-}\to\gamma\chi_{c1}(3872)$ at center-of-mass energies $\sqrt{s}=4.178-4.278$ GeV [22, 23]. Using the $\chi_{c1}(3872)$ signal produced in these data samples, we search for the radiative transition $\chi_{c1}(3872)\to\gamma\psi_{2}(3823)$, where the $\psi_{2}(3823)$ is reconstructed with the cascade decay $\psi_{2}(3823)\to\gamma\chi_{c1}$, $\chi_{c1}\to\gamma J/\psi$, $J/\psi\to\ell^{+}\ell^{-}$ ($\ell=e,\mu$). The branching fraction ratio of this decay relative to the well-established $\chi_{c1}(3872)\to\pi^{+}\pi^{-}J/\psi$ decay, $\mathcal{R}_{\chi_{c1}(3872)}\equiv\mathcal{B}(\chi_{c1}(3872)\to\gamma\psi_{2}(3823),\psi_{2}(3823)\to\gamma\chi_{c1})/\mathcal{B}(\chi_{c1}(3872)\to\pi^{+}\pi^{-}J/\psi)$, is determined in this work.
@@ III Event Selection and result
A four-constraint (4C) kinematic fit is applied to constrain the total four-momentum of the lepton pair and the four photons to that of the colliding beams, to suppress backgrounds and improve the resolution. For events with more than four photons, the combination with the best fit quality corresponding to the minimum fit chi-square, $\chi^{2}_{\rm 4C}$, is retained. The $J/\psi$ is reconstructed by requiring the invariant mass, $M(\ell\ell)$, of the lepton pair to satisfy $|M(\ell\ell)-m(J/\psi)|<30$ MeV/$c^{2}$, where $m(J/\psi)$ is the nominal $J/\psi$ mass. The selection criteria are optimized by maximizing the punzi figure-of-merit $S/(\frac{a}{2}+\sqrt{B})$ [33], where the number of signal events ($S$) is determined with the signal MC sample, the background ($B$) is estimated with the inclusive MC, and the expected statistical significance ($a$) is set to be 3. The dominant background is from the process $e^{+}e^{-}\to\pi^{0}\pi^{0}J/\psi$. After the $J/\psi$ selection, we veto $\pi^{0}$ candidates by requiring that the invariant mass of all photon pairs is more than 15 MeV/$c^{2}$ away from the nominal $\pi^{0}$ mass. After these requirements, a seven-constraint (7C) kinematic fit with an additional three constraints on the masses of $M(\ell\ell)$, $M(\gamma\ell\ell)$, and $M(\gamma\gamma\ell\ell)$ to the nominal masses of $J/\psi$, $\chi_{c1}$, and $\psi_{2}(3823)$, respectively, is applied to distinguish the radiative photon in each cascade decay. The best-fit combination with the minimum chi-square, $\chi^{2}_{\rm 7C}$, is retained; $\chi^{2}_{\rm 7C}<100$ is also required to further suppress the combinatorial backgrounds. One possible peaking background is $\psi_{2}(3823)\to\gamma\chi_{c2},~{}\chi_{c2}\to\gamma J/\psi$, the contribution of which is estimated according to the measurement of the branching fraction ratio of $\psi_{2}(3823)\to\gamma\chi_{c2}$ to $\psi_{2}(3823)\to\gamma\chi_{c1}$ in Ref. [34]. The ratio of the yields of $\psi_{2}(3823)\to\gamma\chi_{c2}$ to $\psi_{2}(3823)\to\gamma\chi_{c1,2}$, is about 1.5%, which is taken into account as a source of systematic uncertainty.
Figure 1 shows the distribution of the invariant mass of the radiative photon and the $\psi_{2}(3823)$, $M(\gamma\psi_{2}(3823))$ for the selected candidates, summed over all the energy points. No signal is observed in the $\chi_{c1}(3872)$ signal region in data. The three events around 3.93 GeV are very unlikely to be from the $\chi_{c2}(2P)$ decays since no $\chi_{c2}(2P)$ signal was observed in its more favourable radiative transition to $\psi(2S)$ [9]. After normalizing the MC samples according to the luminosity and cross section in data, the contributions of the $e^{+}e^{-}\to\pi^{0}\pi^{0}J/\psi$ process and of the other backgrounds, estimated with the inclusive MC sample, are also shown in Fig. 1.
@@
The branching ratio $\mathcal{R}_{\chi_{c1}(3872)}$ is calculated as
@@
where $N_{\rm obs}=0$ is the number of observed events from all data in the $\chi_{c1}(3872)$ signal region [3.855, 3.885] GeV/$c^{2}$ which covers around ±$\pm 3\sigma$ of the signal shape according to the signal MC distributions, $N_{\rm obs}^{\rm sdb}=4$ is the number of the observed events in the $\chi_{c1}(3872)$ sideband region [3.840, 3.855] and [3.885, 3.940] GeV/$c^{2}$; $r$, the background scaling factor from the sideband regions to the signal region, is 0.474 based on the inclusive MC sample (taking into account its systematic uncertainty; see Sec. IV); $N_{\pi^{+}\pi^{-}J/\psi}=80.7\pm 9.0$ is taken from the BESIII measurement [10]; the branching fraction $\mathcal{B}(\chi_{c1}\to\gamma J/\psi)=0.343\pm 0.010$ is quoted from the PDG [2]; $\epsilon_{\gamma\psi_{2}(3823)}$ is the efficiency for the signal process reconstruction, determined with the signal MC sample; and $\epsilon_{\pi^{+}\pi^{-}J/\psi}$ is the efficiency of the process $\chi_{c1}(3872)\to\pi^{+}\pi^{-}J/\psi$ [10]. The efficiency ratio $\epsilon_{\gamma\psi_{2}(3823)}/\epsilon_{\pi^{+}\pi^{-}J/\psi}$ at each energy point is shown in Fig. 2, which is almost independent on the center-of-mass energy. The mean value with the standard deviation, $\epsilon_{\gamma\psi_{2}(3823)}/\epsilon_{\pi^{+}\pi^{-}J/\psi}=0.433\pm 0.004$, is used to calculate the $\mathcal{R}_{\chi_{c1}(3872)}$ value. The upper limit of $\mathcal{R}_{\chi_{c1}(3872)}$ at the 90% confidence level (C.L.) is computed with the TRolke program implemented in the ROOT framework [35] by assuming the background $N_{\rm obs}^{\rm sdb}$ and the denominator of $\mathcal{R}_{\chi_{c1}(3872)}$ follow Poisson and Gaussian distributions, respectively, where the systematic uncertainties discussed in the following section is taken as the standard deviation of the Gaussian function to be considered in the upper limit. We obtain an upper limit of $\mathcal{R}_{\chi_{c1}(3872)}<0.075$ at the 90% C.L.
@@
Systematic uncertainties on $\mathcal{R}_{\chi_{c1}(3872)}$ arise mainly from the estimations of $r$, the possible peaking background of $\psi_{2}(3823)\to\gamma\chi_{c2}\to\gamma\gamma J/\psi$, $N_{\pi^{+}\pi^{-}J/\psi}$, $\epsilon_{\gamma\psi_{2}(3823)}/\epsilon_{\pi^{+}\pi^{-}J/\psi}$, and $\mathcal{B}(\chi_{c1}\to\gamma J/\psi)$. The background scaling factor $r$ is determined from the inclusive MC samples including the process $e^{+}e^{-}\to\pi^{0}\pi^{0}J/\psi$. We use a 1st or 2nd order polynomial function to fit the $M(\gamma\psi_{2}(3823))$ distribution from the inclusive MC samples; the $r$ value is calculated several times using the parameters from the fit and varying them within $1\sigma$. The value $r=0.474$ is chosen from the obtained values since it provides the most conservative upper limit. The contribution of the potential peaking background of $\psi_{2}(3823)\to\gamma\chi_{c2}\to\gamma\gamma J/\psi$ is estimated with the related measurements mentioned previously within one errors, and the result providing the most conservative $\mathcal{R}_{\chi_{c1}(3872)}$ upper limit is retained.